OpenStudy (anonymous):

7 + 6 * ln x - 22.6 = 0

OpenStudy (anonymous):

x= e^3.6

OpenStudy (anonymous):

ln x = 22.6 - 7 / 6 ln x = 2.6 x = e^2.6 = 13.46